OpenStudy (anonymous):

Simplify f(x) = the quantity 3 x-squared minus 4x plus 1 all over 3x minus 1. (4 points)

OpenStudy (anonymous):

\[\frac{ 3x ^{2}-4x+1 }{ 3x-1 }\] This is the expression, i believe it is prime.

OpenStudy (campbell_st):

well factorise the numerator \[\frac{(3x - 1)(x - 1)}{(3x -1)}\] remove the common factor

OpenStudy (anonymous):

Delta of 3x^2-4x+1= 16-4(3)(1)=4 x1=(4-2)/2=1 x2=(4+2)/2=3

OpenStudy (anonymous):

3x-1(x-1)/3x-1=x-1

OpenStudy (anonymous):

Delta? I don't think I've learned that yet. Only in Algebra 2.

OpenStudy (anonymous):

ah ok

Directrix (directrix):

The simplified form would be x - 1 where x is not equal to 1/3. If x = 1/3 the factor (3x-1) becomes zero. Division by 0 is undefined.